[clas_members] Hall B status and shifts

Eugene Pasyuk pasyuk at jlab.org
Fri Jan 13 16:02:41 EST 2023

Dear collaborators,

Here is a brief status update. The configuration change to FT-ON is complete. The detector checkup is ongoing. The solenoid MPS is alive. It was tested with 2400 A with shorted leads.  This would not have happened without the dedicated work of the DC Power group, Magnet group, DSG, Danfysik team, and Hall B technical staff. This was real teamwork. This is very much appreciated. There are a few minor things that remain to be done with MPS. They will be done next week. After that, we would need to do a series of tests with the solenoid. The polarized target also will be cooled down next week as well.
Our original estimate of getting ready for the beam by January 23rd was a bit too optimistic. Now if everything goes well we are looking at January 26-27.
All shifts are canceled until January 25th.    January 26 and up are on standby.

Stay tuned for further updates next week.


From: clas_members <clas_members-bounces at jlab.org> on behalf of Eugene Pasyuk via clas_members <clas_members at jlab.org>
Sent: Tuesday, December 20, 2022 10:43
To: members, clas <clas_members at jlab.org>
Subject: [clas_members] Resumption of RGC and shifts

Dear Collaborators,

First, a brief status update. The configuration change from FT-OUT to FT-ON is ongoing as planned. It is expected to be completed in the first week of January.
As you all know, the RGC run was suspended on November 11 due to the solenoid power supply failure. The root cause of the failure was analyzed and understood:  it was a bug in the power supply firmware. This was confirmed by the vendor, Danfysik. All replacement parts needed for repair have been ordered and have already been received. The magnet group, DC power group, and Hall B engineering are working on repairs. The engineers from Danysik are expected to be at JLab the second week of January. They will bring newly updated firmware. They will review all modifications done to the power supply and will do tests of both power supplies, torus, and solenoid, with new firmware. Once the power supply tests are successful we will proceed with a comprehensive test of the magnets. The test plan is being developed. This will take at least a week.
Under the most optimistic scenario, Hall B may be ready to receive beam on January 23.
With that said, All shifts are canceled until January 23rd. Since there may be a further delay we recommend to those who have shifts assigned for the week of January 23 and require travel consider the possibility to trade their shift with local collaborators. Especially for expert shifts. For the worker shifts, we still have a remote option possible.
I will send updates if anything changes.

Stay tuned and enjoy the holiday break. Hope to see you in the new year in the counting-house.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20230113/6579b4a5/attachment-0001.html>

More information about the clas_members mailing list